xx Creating Constraint Violating Initial Data

Creating Constraint Violating Initial Data

Gabrielle Allen

Abstract

The IDConstraintViolate thorn creates initial data which purposefully violates the constraint equations for testing purposes.

1 Gaussian

Choosing the ADMBase parameter initialxdata to have the value Constraint violating gaussian creates

initial data with the form

where the size and shape of the Gaussian is specified by parameters, =amplitude, =radius and

=sigma.

2 Parameters

11amplitude10Scope: privatex01  REAL

11Description: The amplitude of the waves

11 Range 10 x01 Default: 0.001
11*:*  10No restriction

11

11radius 10Scope: privatex01  REAL

11Description: The radial position of the Gaussian wave

11 Range 10 x01 Default: 0.0
110:*  10Positive

11

11sigma 10Scope: privatex01  REAL

11Description: The sigma for the Gaussian wave

11 Range 10 x01 Default: 0.1
11(0:*  10Positive

11

11initialxdata 10Scope: shared from ADMBASEx01  KEYWORD

11Extends ranges:

11see  [1] below 10Constraint violating Gaussian

11

constraintxviolatingxgaussian

11metricxtype 10Scope: shared from ADMBASEx01KEYWORD

11

3 Interfaces

General Implements: idconstraintviolate

Inherits: admbase grid

4 Schedule This section lists all the variables which are assigned storage by thorn EinsteinInitialData/IDConstraintViolate. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE Scheduled Functions

CCTKxPARAMCHECK x idconstraintviolatexparamchecker x check parameters x

11x11Language:11c
11 x 11Type: 11function

11

ADMBasexInitialData x idconstraintviolatexinitial x set up constraint violating initial data x

11x11Language:11fortran
11 x 11Type: 11function